OpenStudy (anonymous):

The cost to rent a rug-cleaning machine from the supermarket is $23 per day. The supermarket charges a check-out fee of $45 for the rental. Tristanne spent $137 to rent the rug-cleaning machine. If d represents the number of days Tristanne rents the machine, which equation can be used to find the value of d?

OpenStudy (anonymous):

23d=137 23d+45=137 45d=137 45d+23=137

OpenStudy (anonymous):

23d + 45 = 137

OpenStudy (anonymous):

The perimeter of a square is equal to the perimeter of an equilateral triangle. The length of a side of the square is given by x, and the length of a side of the equilateral triangle is given by x=1 . Which equation can be used to find the value of x?

OpenStudy (anonymous):

x=3(x+1) 3x=4(x+1) 4x=3(x+1) x=x+1

OpenStudy (anonymous):

hmm...perimeter of square is 4x and perimeter of equilateral triangle is 3x. so equilateral triangle is 3x + 1. 4x = 3(x + 1)...I think

OpenStudy (anonymous):

you might need to get a second opinion on this one

OpenStudy (anonymous):

it was right

OpenStudy (anonymous):

what a relief....lol :)